A 12 mA steady current is flowing on a copper wire at a time interval of one


minute. Calculate the number of electrons passing on the copper wire

1
Expert's answer
2022-03-16T18:29:29-0400

The net charge passing through the copper wire

"q=It\\
=12*10^{-3}*60=0.72\:\rm C"

The corresponding number of electrons

"N=\frac{q}{e}=\frac{0.72}{1.6*10^{-19}}=4.5*10^{18}"